# Nicholas Michel    Mohamed E. Najd    Ghada Almashaqbeh

### v3

@@ 4.2 System Operation—Transaction Processing
As mentioned before, $\mathsf{ammBoost}$ operates in epochs and rounds. The sidechain committee begins the epoch by retrieving the latest state, i.e., pool token balances, liquidity positions, and user deposits from the mainchain. It then processes all valid sidechain transactions, including swap, mint, burn, and collect (so users send these transactions to the sidechain). These are packaged into meta-blocks such that a meta-block is mined in each round. In the last round of the epoch, this committee produces a summary-block capturing the payouts for participating users, and any changes on liquidity positions, where the updated liquidity pool balances will be computed based on these lists. After that, it invokes the $\mathsf{Sync}$ function in $\mathsf{TokenBank}$ that resembles submitting a sync-transaction to update the AMM state on the mainchain, which is the state of $\mathsf{TokenBank}$.
@@
In this section, we describe how the various transactions are processed and summarized (Figure 5 captures how the sidechain workload is summarized in $\mathsf{ammBoost}$). Before that, we want to point out that although a user will obtain her newly claimed tokens at the end of an epoch, once $\mathsf{TokenBank}$ is synced, they can use these tokens immediately during an epoch for trading. This is because new tokens will be added to the user deposit balance, and those that were used are deducted from deposit. Thus, the latest deposit state reflects the payout a user obtains (which includes refunding any deposit leftover).
@@
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}[\mathsf{in}.\mathsf{type}]\mathrel{-}=\mathsf{tx}.\mathsf{amount}_{\mathsf{in}}$
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}[\mathsf{out}.\mathsf{type}]\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{\mathsf{out}}$
@@
$\mathsf{sum}_{\mathsf{Positions}}[\mathsf{tx}.\mathsf{posId}].\mathsf{amount}_{A}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{A}$
$\mathsf{sum}_{\mathsf{Positions}}[\mathsf{tx}.\mathsf{posId}].\mathsf{amount}_{B}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{B}$
@@
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{A}\mathrel{-}=\mathsf{tx}.\mathsf{amount}_{A}$
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{B}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{B}$
@@
$\mathsf{sum}_{\mathsf{Positions}}[\mathsf{tx}.\mathsf{posId}].\mathsf{amount}_{A}\mathrel{-}=\mathsf{tx}.\mathsf{amount}_{A}$
$\mathsf{sum}_{\mathsf{Positions}}[\mathsf{tx}.\mathsf{posId}].\mathsf{amount}_{B}\mathrel{-}=\mathsf{tx}.\mathsf{amount}_{B}$
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{A}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{A}$
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{B}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{B}$
@@
$\mathsf{sum}_{\mathsf{Positions}}[\mathsf{tx}.\mathsf{posId}].\mathsf{fees}_{A}\mathrel{-}=\mathsf{tx}.\mathsf{amount}_{A}$
$\mathsf{sum}_{\mathsf{Positions}}[\mathsf{tx}.\mathsf{posId}].\mathsf{fees}_{B}\mathrel{-}=\mathsf{tx}.\mathsf{amount}_{B}$
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{A}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{A}$
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{B}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{B}$
@@
In order to execute a swap transaction, a user’s deposit must cover the input token amount. An exact input swap transaction contains: the type and amount of input tokens to be traded, the minimum amount of output tokens the trade will accept (as a protection against slippage), a price limit that the trade should not exceed, and a deadline which is a round number after which the trade becomes invalid if not executed by that time. [^8] The recipient of the traded tokens is by default the issuer of the swap. This can be extended to support stating an explicit recipient that could be different from the issuer. For an exact output swap, the goal is no longer to trade the exact amount of input tokens for the maximum amount of output tokens, but rather to minimize the amount of input tokens required to trade for the desired exact output. As such, the arguments of the function naturally change to reflect that, with the minimum output slippage protection changing to a maximum input slippage protection.
@@
Furthermore, the fees for LPs whose liquidity was used in filling a swap will be computed. To elaborate, when a user submits a swap transaction, they pay a small additional fee, like $0.3\%$ of their transaction’s input or output value. It is paid in the token pair of the pool based on its net liquidity such that the token with the largest amount of net liquidity is used. For example, if a user provides 100 $A$ input tokens in an exact input swap, and token $A$ is the dominant token, then 0.3 $A$ tokens are used for the LP fee, and 99.7 $A$ tokens are used for the swap transaction. These fees are split up proportionally amongst the positions (based on the amount of liquidity they provide) that occupy the price range for which the swap was executed. $\mathsf{ammBoost}$ maintains a per-position fee balance, which is updated on every swap transaction, again using the same logic used by the underlying AMM to compute these fees.
@@
Lastly, recall that a user will not get her actual traded tokens until the end of the epoch. However, she can use these tokens for trading on the sidechain since the sidechain tracks all balances. So basically, the deposit is a tuple of two values; one for each token type. When a swap is executed, the input token amount is deducted from the user’s deposit while the output token amount is added to this deposit, thus allowing the user to use it immediately.
@@
Summary rules. In the summary-block, the committee summarizes all swap transactions as follows: for every client, all her swaps are combined into a single tuple containing: the client public key, the total payout this user should receive. The latter encapsulates both a deduction from her deposit and a refund of any leftover in that deposit. For example, say a user started with a deposit of $(10A,15B)$ and issued one swap during an epoch that traded 5 token $A$ for 10 token $B$. The updated deposit (which will be the payout summary for that user) would be $(5A,25B)$, which represents a payout of 10 $B$ tokens, a deduction of 5 $A$ tokens, and a refund of deposit leftover of $(5A,15B)$. The same logic applies to the rest of the transactions.
@@
Mints. Mint transactions allow the creation of new liquidity positions or modifying existing ones. An LP broadcasts a mint transaction to the sidechain that contains: the lower and upper ticks, representing the price range for which the liquidity is to reside, and the type/amount of the token to be used as liquidity. The mint will be accepted if the issuer LP’s deposit (either the mainchain or sidechain one) can cover the provided liquidity amount.
@@
Processing. This is also processed using the same logic used by the AMM. We resort to a simple approach to track ownership of positions; the sidechain committee generates a unique identifier (e.g., the hash of the mint transaction and the LP’s public key) for a new position, and the owner is the public key of the issuer LP. An existing position will receive an increase in its balance (or any other modifications on its price range) after verifying that the transaction issuer is indeed the rightful position owner. Mint transactions are initially invoked with a desired amount of token $A$ and token $B$ as input. The underlying AMM algorithms compute the maximum amount of liquidity (based off the input tokens) that the pool can take in at the current moment from both token types. These values represent the share of the pool liquidity now owned by the newly minted or modified position. The committee then deducts the provided liquidity amount (from both token types) from the corresponding LP’s deposit balance.
@@
Summary rules. All mint transactions are summarized as a list of liquidity positions with each position consisting of a tuple containing: the position identifier, the public key of the owner, the total amount of liquidity provided (or net change) for each token type under this position, and total amount of accrued fees. Note that the payin/payout of the LP is also updated when summarizing mint transactions; all provided liquidity token amounts are deducted from their deposits as shown in Figure 5.
@@
Burns. A burn transaction allows a partial or complete liquidity withdrawal of a position. It is issued by an LP and contains: the position ID, the tick price limits, and the desired amounts of tokens $A$ and $B$ to be burned, and sent to the sidechain.
@@
Processing. Processing a burn transaction boils down to determining if the issuer LP owns the position, then calculating the amount of liquidity this LP owns in a share, and converting that amount of liquidity into an amount of both tokens managed by the pool (using the original logic of the AMM). This would lead to updating the position range (upper and lower price ticks), or a deletion of the position if all its associated liquidity is withdrawn. If a deleted position has fees owed to it, the owner LP will receive these fees as part of her total payout computed at the end of the epoch, i.e., will be added to their deposit balance.
@@
Collects. Collects allow LPs to collect the fees earned by their liquidity positions. An LP broadcasts a collect transaction containing the identifier of the position and fee amount to be collected.
@@
Flashes. Flash transactions allow users to request short-term loans within the duration of one mainchain block. These are the only transaction type that $\mathsf{ammBoost}$ does not offload to the sidechain; the delay in paying out the actual tokens (which happens at the end of an epoch) limits the intended use of flash loans that span a very short period. As such, in $\mathsf{ammBoost}$, flash transactions happen on the mainchain as in the original AMM architecture. Since flash loans take place in a singular block, they do not impact the pool balances; the amount of loaned tokens should be returned within one block period or the loan will be inverted. As a result, they do not invalidate any of the transactions processed on the sidechain based off the balance snapshot taken at the beginning of an epoch.
Remark 1
In terms of user experience, clients and LPs should be connected to both the mainchain and the sidechain, and their wallets should issue transactions to the destination chain based on the transaction type. For example, deposits should go the mainchain, while swaps/mints/burns/collects should go to the sidechain. Another difference is related to receiving the actual tokens, which are delayed until the end of the epoch in $\mathsf{ammBoost}$. Since a user can use these immediately for trades within an epoch, the delayed payout has no impact. However, if a user wants to use these tokens on a different AMM, or wants to trade these with other token types to participate in another pool managing different token pairs, then they have to wait until the epoch. Still, overall, the delay is only one epoch.
4.3 System Operation—Chain Management
@@
Syncing $\mathsf{TokenBank}$. The sidechain committee leader, after producing a summary-block containing all the summaries detailed earlier, calls the $\mathsf{Sync}$ function in $\mathsf{TokenBank}$ that resembles a sync-transaction submission. The inputs to this function call include: the list of payouts for all clients and LPs, and the list of liquidity positions with their updated information.
@@
Authentication. $\mathsf{TokenBank}$ must ensure that the the $\mathsf{Sync}$ function invocation is issued by the rightful sidechain committee. We use a modified idea of quorum certificates (QC) [54, 72] combined with threshold signatures. In detail, to authenticate the $\mathsf{Sync}$ call for epoch $e+1$, the election of committee $e+1$ must happen during epoch $e$. Then, this committee runs a distributed key generation (DKG) [35] to generate a public verification key $\mathsf{vk}_{c}$ for the committee and secret shares of the signing key (one share per member) with a threshold of $22f+$ 2.999A committee size is $23f+$ 2 and $f$ is the maximum number of faulty nodes as in PBFT protocols, $22f+$ 2 votes are needed to reach an agreement. This committee initiates an agreement on $\mathsf{vk}_{c}$, and then sends the agreement output to committee $e$ along with proofs of election of each member who participated in the agreement. [^10] In our implementation, this election proof is the output of the verifiable random function (VRF) used in the election mechanism. Committee $e$ verifies the election proofs, and then verifies that there is an agreement on $\mathsf{vk}_{c}$. If everything is correct, committee $e$ records $\mathsf{vk}_{c}$ on $\mathsf{TokenBank}$ by adding that to the $\mathsf{Sync}$ function call inputs they submit at the end of epoch $e$. During epoch $e+1$, committee $e+1$ runs an agreement over the $\mathsf{Sync}$ function call inputs and signs using their signing key shares, which result in one signature over these inputs. The leader then invokes $\mathsf{Sync}$ with the inputs and this signature. In turn, $\mathsf{TokenBank}$ verifies the signature using the recorded $\mathsf{vk}_{c}$ before accepting the summaries. By the security of the threshold signature scheme, this signature will be valid only if at least $22f+$ 2 committee members has signed.
@@
Processing. If successfully verified, $\mathsf{TokenBank}$ processes the $\mathsf{Sync}$ function call by updating the list of positions based on the summaries, i.e., delete fully withdrawn positions, create new positions, or adjust existing ones, Then, it updates the pool balance based on the reported payouts and updated position list. Lastly, it dispenses the payouts to the referenced clients/LPs.
@@
Sidechain pruning. $\mathsf{ammBoost}$ uses the block suppression technique from chainBoost. Once the transaction encapsulating the $\mathsf{Sync}$ function call is confirmed on the mainchain, all meta-blocks associated to this transaction will be pruned. The summary-blocks, as mentioned before, are permanent and represent checkpoints of the sidechain state in each epoch. So they can be used to verify the state of the AMM reflected by $\mathsf{TokenBank}$ state variables.
@@
Handling interruptions. We identify the scenarios that can lead to operation interruption in $\mathsf{ammBoost}$ and how to recover from them. Recall that the sidechain committees use a PBFT-based consensus that assumes up to $f$ of the elected miners can be malicious. [^11] This is valid under a committee size that guarantees satisfying this condition with overwhelming probability, where we adopt the committee size analysis from [57]. Thus, interruptions that could happen result from having a malicious or unresponsive leader. This leader may either propose an invalid meta/summary-blocks or invalid function call to $\mathsf{Sync}$, or not initiate the agreement in the first place. Another interruption could result from rollbacks on the mainchain. That is, when the mainchain miners switch their canonical chain to the one satisfying a particular fork resolution criteria (i.e. the longest branch, or the heaviest one), causing the most recent blocks to be abandoned. This is an issue if the abandoned blocks contain $\mathsf{Sync}$ transactions.
@@
Detection and recovery from these interruptions are done as in chainBoost [57], which we review briefly here. A leader that proposes an invalid block or $\mathsf{Sync}$ call can be easily detected by the committee when verifying this proposal. Once detected, the view-change technique [41] is used to elect a new leader. In the case of an unresponsive leader, if no agreement is initiated within a timeout period, a leader change is triggered. As for a leader that proposes invalid $\mathsf{Sync}$ inputs, a leader-change will not help since this happens at the end of an epoch when it is time for the new committee to take over. Thus, this case, and the rollback interruption, are addressed using the mass-syncing technique. The new committee issues a $\mathsf{Sync}$ call covering the summaries they produced in their epoch and those produced earlier in the impacted epochs.
Remark 2
@@
A sidechain operates as a regular blockchain, thus any transactions that have not been processed in an epoch will be carried over to the epoch after. All sidechain miners receive transactions destined to the sidechain, but only the elected committee mines meta and summary blocks. Thus, when a new meta-block is mined, the committee and all other sidechain miners remove all published transactions in that block from their queues.
4.4 Security

### v4

@@ IV-B System Operation—Transaction Processing
$\mathsf{ammBoost}$ operates in epochs and rounds. The sidechain committee begins the epoch by retrieving the latest state, i.e., pool token balances, liquidity positions, and user deposits from the mainchain. It then processes all valid sidechain transactions, including swap, mint, burn, and collect (so users send these transactions to the sidechain). These are packaged into meta-blocks such that a meta-block is mined in each round. In the last round of the epoch, this committee produces a summary-block capturing the payouts for participating users, and any changes on liquidity positions, where the updated liquidity pool balances will be computed based on these lists. After that, it invokes the $\mathsf{Sync}$ function in $\mathsf{TokenBank}$ that resembles submitting a sync-transaction to update the AMM state on the mainchain, which is the state of $\mathsf{TokenBank}$.
@@
In this section, we describe how the various transactions are processed and summarized (Figure 4 captures how the sidechain workload is summarized in $\mathsf{ammBoost}$). Before that, we want to point out that although a user will obtain her newly claimed tokens at the end of an epoch, still they can use these tokens immediately during an epoch for trading. This is because new tokens will be added to the user deposit balance, and those that were used are deducted from deposit. Thus, the latest deposit state includes the newly accrued tokens, and it can be used for any trading activity by the user.
@@
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amnt}[\mathsf{in}.\mathsf{type}]\mathrel{-}=\mathsf{tx}.\mathsf{amnt}_{\mathsf{in}}$
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amnt}[\mathsf{out}.\mathsf{type}]\mathrel{+}=\mathsf{tx}.\mathsf{amnt}_{\mathsf{out}}$
@@
$\mathsf{sum}_{\mathsf{Positions}}[\mathsf{tx}.\mathsf{posId}].\mathsf{amnt}_{A}\mathrel{+}=\mathsf{tx}.\mathsf{amnt}_{A}$
$\mathsf{sum}_{\mathsf{Positions}}[\mathsf{tx}.\mathsf{posId}].\mathsf{amnt}_{B}\mathrel{+}=\mathsf{tx}.\mathsf{amnt}_{B}$
@@
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amnt}_{A}\mathrel{-}=\mathsf{tx}.\mathsf{amnt}_{A}$
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amnt}_{B}\mathrel{+}=\mathsf{tx}.\mathsf{amnt}_{B}$
@@
$\mathsf{sum}_{\mathsf{Positions}}[\mathsf{tx}.\mathsf{posId}].\mathsf{amnt}_{A}\mathrel{-}=\mathsf{tx}.\mathsf{amnt}_{A}$
$\mathsf{sum}_{\mathsf{Positions}}[\mathsf{tx}.\mathsf{posId}].\mathsf{amnt}_{B}\mathrel{-}=\mathsf{tx}.\mathsf{amnt}_{B}$
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amnt}_{A}\mathrel{+}=\mathsf{tx}.\mathsf{amnt}_{A}$
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amnt}_{B}\mathrel{+}=\mathsf{tx}.\mathsf{amnt}_{B}$
@@
$\mathsf{sum}_{\mathsf{Positions}}[\mathsf{tx}.\mathsf{posId}].\mathsf{fees}_{A}\mathrel{-}=\mathsf{tx}.\mathsf{amnt}_{A}$
$\mathsf{sum}_{\mathsf{Positions}}[\mathsf{tx}.\mathsf{posId}].\mathsf{fees}_{B}\mathrel{-}=\mathsf{tx}.\mathsf{amnt}_{B}$
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amnt}_{A}\mathrel{+}=\mathsf{tx}.\mathsf{amnt}_{A}$
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amnt}_{B}\mathrel{+}=\mathsf{tx}.\mathsf{amnt}_{B}$
@@
In order to execute a swap transaction, a user’s deposit must cover the input token amount. An exact input swap transaction contains: the type and amount of input tokens to be traded, the minimum amount of output tokens the trade will accept (as a protection against slippage), a price limit that the trade should not exceed, and a deadline which is a round number after which the trade becomes invalid if not executed by that time. The recipient of the traded tokens is the issuer of the swap (this can be extended to support stating an explicit recipient other than the issuer). For an exact output swap, the goal is no longer to trade the exact amount of input tokens for the maximum amount of output tokens, but rather to minimize the amount of input tokens required to trade for the desired exact output. As such, the arguments of the function naturally change to reflect that, with the minimum output slippage protection changing to a maximum input slippage protection.
@@
Furthermore, the fees for LPs whose liquidity was used in filling a swap will be computed. To elaborate, when a user submits a swap transaction, they pay a small additional fee in the token pair of the pool. These fees are split proportionally among the positions (based on the amount of liquidity they provide) that occupy the price range for which the swap was executed. $\mathsf{ammBoost}$ maintains a per-position fee balance, which is updated after every swap, again using the same logic used by the underlying AMM to compute these fees.
@@
Lastly, recall that a user will not get her actual traded tokens until the end of the epoch. However, she can use these tokens for trading on the sidechain since the sidechain tracks all balances. So basically, the deposit is a tuple of two values; one for each token type. When a swap is executed, the input token amount is deducted from the user’s deposit while the output token amount is added to this deposit, thus allowing the user to use the newly accrued tokens immediately.
@@
Summary rules. The committee summarizes swaps as follows: for every client, all their swaps are combined into a single tuple containing the client public key and the total payout they should receive. The latter encapsulates both a deduction from their deposit and a refund of any deposit leftover. For example, say a client started with a deposit of $(10A,15B)$ and issued one swap during an epoch that traded 5 token $A$ for 10 token $B$. The updated deposit would be $(5A,25B)$, which represents a payout of 10 $B$, a deduction of 5 $A$, and a refund of deposit leftover of $(5A,15B)$.
@@
Mints. Mint transactions allow the creation of new liquidity positions or modifying existing ones. An LP broadcasts a mint transaction to the sidechain that contains: the lower and upper ticks, representing the price range for which the liquidity can be used, and the type/amount of the token to be used as liquidity. The mint will be accepted if the issuer LP’s deposit can cover the provided liquidity amount.
@@
Processing. This is also processed using the same logic used by the AMM. We employ a simple approach to track ownership of positions; the sidechain committee generates a unique identifier (e.g., the hash of the mint transaction and the LP’s public key) for a new position, and the owner is the public key of the issuer LP. An existing position will receive an increase in its balance (or any other modifications on its price range) after verifying that the transaction issuer is indeed the rightful position owner. Mint transactions are initially invoked with a desired amount of token $A$ and token $B$ as input. The underlying AMM algorithms compute the maximum amount of liquidity (based off the input tokens) that the pool can take in at the current moment from both token types. These values represent the share of the pool liquidity now owned by the newly minted or modified position. The committee then deducts the provided liquidity amount (from both token types) from the corresponding LP’s deposit balance.
@@
Summary rules. All mints are summarized as a list of liquidity positions with each position consisting of a tuple containing: the position identifier, the public key of the owner, the total amount of liquidity provided (or net change) for each token type, and total amount of accrued fees. Note that the payin/payout of the LP is also updated when summarizing mint transactions; all provided liquidity token amounts are deducted from their deposits as shown in Figure 4.
@@
Burns. A burn transaction allows a partial or complete liquidity withdrawal of a position. It is issued by an LP and contains: the position ID, the tick price limits, and the desired amounts of tokens $A$ and $B$ to be burned.
@@
Processing. Processing a burn transaction boils down to determining if the issuer LP owns the position, then calculating the amount of liquidity this LP owns in a share, and converting that amount of liquidity into an amount of both tokens managed by the pool (using the original logic of the AMM). This would lead to updating the position range (upper and lower price ticks), or a deletion of the position if all its associated liquidity is withdrawn. If a deleted position has fees owed to it, the owner LP will receive these fees as part of her total payout computed at the end of the epoch, i.e., will be added to her deposit balance.
@@
Collects. Collects allow LPs to collect the fees earned by their liquidity positions. An LP broadcasts a collect transaction containing the position identifier and fee amount to be collected.
@@
Flashes. Flash transactions allow users to request short-term loans within the duration of one mainchain block. The goal is to exploit arbitrage opportunities across trading platforms (e.g., different AMMs or other exchanges), thus achieving quick profits due to token price discrepancies. As a result, different from swaps/mints/collects, flash loans are not an internal AMM operation, but rather a short-term lending service that some AMMs offer, and it needs a transfer of actual tokens to the loanee rather than tracking internal token balances.
Flash loans are the only transaction type that $\mathsf{ammBoost}$ does not offload to the sidechain; the delay in paying out the actual tokens (which happens at the end of an epoch) limits the intended use of flash loans that span a very short period. As such, flash transactions happen on the mainchain as in the original AMM architecture. Since flash loans take place in a singular block, they do not impact the pool balances; the loaned tokens must be returned within one block period or the loan will be inverted. As a result, they do not invalidate any of the transactions processed on the sidechain based off the balance snapshot taken at the beginning of an epoch.
Remark 1.
In terms of user experience, clients and LPs should be connected to the side and main chains, with their wallets issuing transactions each chain based on the transaction type. Also, receiving actual tokens is delayed until the end of the epoch. Since a user can use these immediately for trades within an epoch, the delayed payout has no impact. However, if a user wants to use these tokens on a different AMM, for example, then they have to wait until the epoch ends. Nonetheless, this delay is still way shorter than operating the AMM solely on the mainchain (i.e., without $\mathsf{ammBoost}$) due to congestion/low throughput issues of blockchains as discussed in the introduction and as we show in Section VI.
IV-C System Operation—Chain Management
@@
Syncing $\mathsf{TokenBank}$. The sidechain committee leader, after producing a summary-block for an epoch, calls the $\mathsf{Sync}$ function in $\mathsf{TokenBank}$ that resembles a sync-transaction submission. The inputs to this function call include: the list of payouts for all clients and LPs, and the list of liquidity positions with their updated information.
@@
Authentication. $\mathsf{TokenBank}$ must ensure that a $\mathsf{Sync}$ invocation is issued by the rightful sidechain committee. We use the idea of quorum certificates (QC) [82, 83] combined with threshold signatures, which we refer to as TSQC. To authenticate the $\mathsf{Sync}$ call for epoch $e+1$, the election of committee $e+1$ must happen during epoch $e$. Then, this committee runs a distributed key generation (DKG) protocol [84] to generate a public verification key $\mathsf{vk}_{c}$ for the committee and secret shares of the signing key (one share per member) with a threshold of $22f+$ 2. This committee initiates an agreement on $\mathsf{vk}_{c}$, and then sends the agreement output to committee $e$ along with proofs of election of each member who participated in the agreement. [^4] In our implementation, this election proof is the output of the verifiable random function (VRF) used in the election mechanism. Committee $e$ verifies the election proofs, and then verifies that there is an agreement on $\mathsf{vk}_{c}$. It then records $\mathsf{vk}_{c}$ on $\mathsf{TokenBank}$ by adding that to the $\mathsf{Sync}$ function call inputs they submit at the end of epoch $e$. During epoch $e+1$, committee $e+1$ runs an agreement over the $\mathsf{Sync}$ function call inputs and signs using their signing key shares, which result in one signature over these inputs. The leader then invokes $\mathsf{Sync}$ with the inputs and this signature. In turn, $\mathsf{TokenBank}$ verifies the signature using the recorded $\mathsf{vk}_{c}$. By the security of the threshold signature scheme, this signature will be valid only if at least $22f+$ 2 committee members has signed.
@@
Processing. If successfully verified, $\mathsf{TokenBank}$ processes the $\mathsf{Sync}$ function call by updating the list of positions based on the summaries, i.e., delete fully withdrawn positions, create new positions, or adjust existing ones, Then, it updates the pool balance based on the reported payouts/updated positions. Lastly, it dispenses the payouts to the referenced clients/LPs.
@@
Sidechain pruning. $\mathsf{ammBoost}$ uses the block suppression technique from chainBoost. Once the transaction encapsulating the $\mathsf{Sync}$ call is confirmed on the mainchain, all meta-blocks associated to this transaction will be pruned. Summary-blocks, as mentioned before, are permanent and represent checkpoints for sidechain state in each epoch. So they can be used to verify the state of the AMM reflected by $\mathsf{TokenBank}$ state variables.
@@
Handling interruptions. We identify the scenarios that can lead to operation interruption in $\mathsf{ammBoost}$ and how to recover from them. Recall that sidechain committees use a secure PBFT-based consensus that assumes up to $f$ of the elected miners can be malicious. This is valid under a committee size that guarantees satisfying this condition with overwhelming probability, where we adopt the committee size lower bound from [67]. Thus, a committee will not agree on invalid blocks, and the only interruptions a committee might have are due to having a malicious or unresponsive leader. This leader may either propose an invalid meta/summary-blocks or invalid function call to $\mathsf{Sync}$, or not initiate the agreement in the first place. Another interruption could result from rollbacks on the mainchain. That is, when the mainchain miners switch their canonical chain to the one satisfying a particular fork resolution criteria (i.e., the longest branch, or the heaviest one), causing the most recent blocks to be abandoned. This is an issue if the abandoned blocks contain $\mathsf{Sync}$ transactions.
@@
Detection and recovery from these interruptions are done as in chainBoost [67]. Briefly, a leader that proposes an invalid block or $\mathsf{Sync}$ call will be easily detected by the committee when verifying this proposal. Once detected, view-change [85] triggers electing a new leader. For an unresponsive leader, if no agreement is initiated within a timeout period, view-change is invoked. For a leader that proposes invalid $\mathsf{Sync}$ inputs, view-change will not help since this happens at the epoch end when it is time for the new committee to take over. Hence, this case, and the rollback interruption, are addressed using mass-syncing. The new committee issues a $\mathsf{Sync}$ call covering the summaries they produced in their epoch and those produced earlier in the impacted epochs.
Remark 2.
@@
A sidechain operates as a regular blockchain, thus any transactions that have not been processed in an epoch will be carried over to the epoch after. All sidechain miners receive transactions destined to the sidechain, but only the elected committee mines meta and summary blocks. Thus, when a new meta-block is mined, the committee and all other sidechain miners remove all published transactions from their queues.
IV-D Security